# Manipulation Facing Threats: Evaluating Physical Vulnerabilities in End-to-End Vision Language Action Models

### v2

@@ Abstract
Recently, driven by advancements in Multimodal Large Language Models (MLLMs), Vision Language Action Models (VLAMs) are being proposed to achieve better performance in open-vocabulary scenarios for robotic manipulation tasks. Since manipulation tasks involve direct interaction with the physical world, ensuring robustness and safety during the execution of this task is always a very critical issue. In this paper, by synthesizing current safety research on MLLMs and the specific application scenarios of the manipulation task in the physical world, we comprehensively evaluate VLAMs in the face of potential physical threats. Specifically, we propose the Physical Vulnerability Evaluating Pipeline (PVEP) that can incorporate as many visual modal physical threats as possible for evaluating the physical robustness of VLAMs. The physical threats in PVEP specifically include Out-of-Distribution, Typography-based Visual Prompt, and Adversarial Patch Attacks. By comparing the performance fluctuations of VLAMs before and after being attacked, we provide generalizable Analyses of how VLAMs respond to different physical security threats. Our project page is in this link
@@ I Introduction
As a task with widespread applications in real-life and industrial manufacturing [negrello2020hands, ghodsian2023mobile, li2022robotic, davidson2020robotic, wu2023tidybot, yang2023transferring, wu2024new, yoshikawa2023large], the continuous performance improvements of robotic arm manipulation systems in recent years have been driven by the development of various Artificial Intelligence (AI) algorithms. Previously, research on AI-driven manipulation systems predominantly focused on training-from-scratch imitation learning methods, such as behavior cloning [florence2022implicit, wang2023identifying] and diffusion policy [ma2024hierarchical, xian2023chaineddiffuser, ze20243d]. However, with the emergence of Large Language Models (LLMs) [alayrac2022flamingo, touvron2023llama] as well as Multimodal Large Language Models (MLLMs) [zhu2023minigpt, liu2023llava, liu2023improvedllava], LLMs/MLLMs-driven manipulation systems have also been introduced.
@@
Figure 1: Performance degradation and time delay of LLaRA and OpenVLA due to physical attacks (×$\times 35$ is for better illustration).
@@
Compared to imitation learning models trained on single tasks, LLMs/MLLMs-driven manipulation systems could gain better performance in open-vocabulary scenarios because of the powerful informative capacity of large models. Among these, systems that directly leverage commercial closed-source API to enhance manipulation task performance in open-vocabulary situations have been applied in various areas [wu2023tidybot, yang2023transferring, wu2024new, yoshikawa2023large]. However, due to the inconvenience of using commercial closed-source models and the challenges of specific deployment, as well as the goal of achieving real AI democratization, it is crucial to propose the LLMs/MLLMs-driven manipulation systems that allows all researchers to access complete information. Based on this, manipulation systems leveraging open-source LLMs have also been proposed. However, unlike the simultaneous improvement in visual perception and instruction semantic understanding brought by incorporating commercial large model API, open-source LLMs-driven systems still rely on traditional vision models, such as YOLO [redmon2016you] and Mask R-CNN [he2017mask], for the visual perception module to acquire visual modality information. Traditional vision models are typically designed for specific, single-vision tasks, which result in a lack of generality and scalability when applied to different zero-shot tasks. During the training process, these models learn from manually annotated, simple vision-label information pairs, which limits their ability to develop a more comprehensive and complex understanding of visual information when combined with rich language descriptions. This would incur the following limitations for LLMs-driven manipulation systems in open-vocabulary scenarios: (1) inability to perceive and handle unseen or zero-shot novel objects and environments; (2) inability to understand complex semantic instructions, which prevents the system from responding to more sophisticated tasks. To overcome the aforementioned limits, building upon the design principles of MLLMs, the end-to-end Vision-Language-Action Models (VLAMs) have been proposed.
@@
As for the physical security threats to AI-driven manipulation systems, [tsurumine2022goal] and [chen2024diffusion] investigate the potential safety issues associated with using imitation learning. [wu2024safety, li2024mmro, zhang2024badrobot] explore how commercial API and open-source LLMs-driven manipulation systems experience performance degradation due to hallucination and jailbreak issues. As for the MLLMs-driven end-to-end VLAMs, no related physical security validation work has been introduced so far. Therefore, this paper conducts a comprehensive robustness and safety evaluation of open-source VLAMs, represented by LLaRA [li2024llara] and OpenVLA [kim2024openvla]. For the specific evaluation approaches that may arise when applying VLAMs under fully open-vocabulary scenarios within the physical world. We do not include attack methods targeting the language instructions in our threat considerations. The specific reason is that in physical attack scenarios, VLAMs mainly face relatively unchanged visual modality information, along with various semantic and linguistic template language instructions to process. Therefore, compared to constantly modifying language modality input, attacks targeting the visual modality input would have a more profound impact on the physical world. Based on this, we propose the Physical Vulnerability Evaluation Pipeline (PVEP) to assess the performance variations of VLAMs when confronting visual modality safety threats. PVEP includes most currently possible visual attacks in the physical world, such as Out-of-Distribution (OOD), Typography-based Visual Prompt (VP), and Adversarial Patch (AdvP) Attacks, to the best of our knowledge. Measured by the failure rate and the number of timesteps to complete tasks, Figure 1 shows the largest performance degradation and corresponding time delay of VLAMs before (Clean) and after suffering various physical attacks (OOD, VP, AdvP). Furthermore, through comprehensive evaluations in PVEP, we present four generalizable Analyses for VLAMs in open-vocabulary scenarios:
@@ II Background
LLMs/MLLMs-driven Manipulation systems: The recent emergence of commercial closed-source large-model APIs, represented by GPT-4 [achiam2023gpt] and Claude [claude], marks a significant step forward in the research toward Artificial General Intelligence (AGI). In contrast to closed-source models, numerous open-source LLMs and MLLMs are proposed to promote the democratization of AI development. For open-source LLMs, Flamingo [alayrac2022flamingo] LLama [touvron2023llama], Vicuna [kassem2024alpaca] and others [yao2023llm, rawte2023survey, yin2023survey] obtain good performance in various zero-shot natural language processing tasks. Open-source MLLMs, such as LLaVA, MiniGPT-4 and others  [zhu2023minigpt, liu2023llava, liu2023improvedllava, karamcheti2024prismatic], could be obtained through fusing open-source pre-trained vision encoders [radford2021learning] with different open-source LLMs. Based on this progress, there are lots of commercial API-driven manipulation systems have been adopted widespread applications in different areas, such as home-services robot [wu2023tidybot, yang2023transferring], agriculture robot [wu2024new], scientific experiments robot [yoshikawa2023large]. Then, referring to the advancement of open-source LLMs, some LLMs-driven manipulation systems, such as VIMA and others [jiang2022vima, nair2022r3m], are also proposed. Additionally, about MLLMs-driven manipulation systems, a promising trend has emerged in the form of Vision-Language-Action models (VLAMs), which involve fine-tuning large pre-trained MLLMs for robot action prediction. These models are characterized by their fusion of robot control actions directly into MLLMs backbones. RT-2 [brohan2023rt] and PaLM [anil2023palm, driess2023palm] have garnered significant attention due to their claims of achieving promising performance across various manipulation tasks. However, both of them are closed-source. Therefore, LLaRA [li2024llara] and OpenVLA [kim2024openvla] as the totally open-source VLAMs are proposed. Since open-source models can be continuously evolved due to their easy accessibility, these two VLAMs may potentially revolutionize autonomous systems and human-robot interaction paradigms in the future.
@@
Safety Concerns: Different types of physical visual attacks could significantly impact the performance of various AI models. OOD attacks [lee2018simple, besnier2021triggering, herrmann2022pyramid, wang2024llms] can consistently affect the stability of AI models. For the recently emerged LLMs and MLLMs, there are also issues such as jailbreak [wei2024jailbroken, huang2023catastrophic, wang2024llms, xu2024defending] and hallucination [yao2023llm, rawte2023survey, tonmoy2024comprehensive, xu2024hallucination] that can undermine the reliability of the final language output. Besides, for MLLMs, typography-based visual prompts [azuma2023defense, cheng2024typographic] could distract the semantics of final language output by adding simple pixel-level text to visual modality input. In addition, various kinds of AI models, including LLMs and MLLMs, as victim models, have been certified to possess adversarial vulnerabilities [bai2021recent, mao2022towards, kumar2023certifying, wang2024stop]. The output of AI models can be altered by adversarial perturbations, leading to either deviation from the correct output (untargeted) or convergence towards a predefined incorrect output (targeted). Furthermore, the generating process of such perturbations could be classified into black and white box, depending on the amount of information attackers possess about the victim model. About the particular vulnerability of AI-driven manipulation systems, [tsurumine2022goal] and [chen2024diffusion] investigate the potential safety issues of using diffusion policy during executing manipulation tasks. [zhang2024badrobot] explores the jailbreak threats in commercial API-driven manipulation systems. And [li2024mmro] introduces the MMRo benchmark, designed to evaluate API-driven systems across four key areas: perception, task planning, visual reasoning, and safety. [wu2024safety] evaluates safety challenges of open-source LLMs-driven manipulation systems in vision-language modality. However, there is no physical threat evaluation work targeting open-source VLAMs so far.
@@ III-B Threat Models
Additionally, it is crucial to analyze the impact of specific types of input image-language modality pairs on the adversarial transferability performance of VLAMs. Algorihtm 1 presents the execution process of the adversarial transfer attack from MLLMs to VLAMs. For black-box Attackers, compared to robotic vision-instruction language pairs $\{x_{rv},t_{vi}\}$ for VLAMs, which typically requires simulation or real-world scene recording to obtain, the general VQA image-prompt pairs $\{x_{gi},t_{gp}\}$ of MLLMs are often more easily available as the online dataset, such as DAQUAR [malinowski2014multi], TallyQA[acharya2019tallyqa], A-OKVQA [schwenk2022okvqa] and others.
@@
4:$\delta\leftarrow\min_{L}(f_{m}({\theta}_{m},x_{i}\odot(1-m)+\delta\odot m,t_{j}),y_{L})$triangleright $\{x_{i},t_{j}\}$ could be $\{x_{gi},t_{gp}\}$ or $\{x_{rv},t_{ri}\}$
5:$y_{t}\leftarrow f_{v}({\theta}_{v},x_{rv}\odot(1-m)+\delta\odot m,t_{ri})$triangleright Apply $\delta$ generated from MLLMs to VLAMs
@@ IV Experiments
Failure Rate (%) Clean Blur GN BC(B) BC(D) TW1 TW2 TW3 TW4 TN1 TN2 TN3 BB RBB GB WB
@@ IV-A Models and Simulators
For VLAMs, we adopt LLaRA [li2024llara] and OpenVLA [kim2024openvla] as victim models. The LLaRA and OpenVLA are evaluated on the VIMA [jiang2022vima] and SimplerEnv [li2024evaluating] simulator respectively.
@@ IV-B3 Adversarial Patch Attack
In particular, for the BB attack, we utilize minimal information about the victim VLAMs. We employe only the victim VLAMs’ base model (pre-fine-tuning) as a surrogate model. For instance, for the victim model LLaRA, we use its base model LLaVA as the surrogate model. The adversarial patch was trained using 5000 images from ImageNet and 200 general Visual Question Answering (VQA) prompts selected from [luo2024image]. For the RBB attack, based on BB, we provide additional data information for training the adversarial patch. We collect an additional 2000 Images and 200 prompts of robotics manipulation scenarios when using victim VLAMs for inference. For the GB attack, we use the victim VLAMs and the general VQA images/prompts in BB for adversarial patch training. As for the WB attack, we directly use the victim VLAMs and all available images and prompts from both BB and RBB. By systematically comparing these four levels of adversarial patch attacks, we can assess the minimum level of model and data information required to generate an effective adversarial patch against VLAMs.
@@
Failure Rate (%) Clean Blur GN BC(B) BC(D) TW1 TW2 TW3 TW4 TN1 TN2 TN3 BB RBB WB
@@ V Conclusion and Limitation
Limitation: Since we are evaluating the most cutting-edge open-source VLAMs recently proposed, the official documentation regarding their deployment details on real-world robotic arms is still being updated. Additionally, OpenVLA does not directly provide a simulation validation method, and the current mainstream approach is to use SimplerEnv [li2024evaluating]. However, this combination involves a significant workload and still shows suboptimal performance on certain tasks, which [simpler1, simpler2] is actively working to improve. We will continue to update the results and further conduct validation on physical systems. Also, when new VLAMs are proposed, we will utilize PVEP for further validation.
[…]

### v3

@@ Abstract
Recently, driven by advancements in Multimodal Large Language Models (MLLMs), Vision Language Action Models (VLAMs) are being proposed to achieve better performance in open-vocabulary scenarios for robotic manipulation tasks. Since manipulation tasks involve direct interaction with the physical world, ensuring robustness and safety during the execution of this task is always a very critical issue. In this paper, by synthesizing current safety research on MLLMs and the specific application scenarios of the manipulation task in the physical world, we comprehensively evaluate VLAMs in the face of potential physical threats. Specifically, we propose the Physical Vulnerability Evaluating Pipeline (PVEP) that can incorporate as many visual modal physical threats as possible for evaluating the physical robustness of VLAMs. The physical threats in PVEP specifically include Out-of-Distribution, Typography-based Visual Prompt, and Adversarial Patch Attacks. By comparing the performance fluctuations of VLAMs before and after being attacked, we provide generalizable Analyses of how VLAMs respond to different physical threats.
@@ I Introduction
As a task with widespread applications in real-life and industrial manufacturing [1, 2, 3, 4, 5, 6, 7, 8], the continuous performance improvements of robotic arm manipulation systems in recent years have been driven by the development of various Artificial Intelligence (AI) algorithms. Previously, research on AI-driven manipulation systems predominantly focused on training-from-scratch imitation learning methods, such as behavior cloning [9, 10] and diffusion policy [11, 12, 13]. However, with the emergence of Large Language Models (LLMs) [14, 15] as well as Multimodal Large Language Models (MLLMs) [16, 17, 18], LLMs/MLLMs-driven manipulation systems have also been introduced.
@@
Figure 1: Performance degradation and time delay of LLaRA and OpenVLA due to physical attacks ($\times 35$ is for better illustration).
@@
Compared to imitation learning models trained on single tasks, LLMs/MLLMs-driven manipulation systems could gain better performance in open-vocabulary scenarios because of the powerful informative capacity of large models. Among these, systems that directly leverage commercial closed-source API to enhance manipulation task performance in open-vocabulary situations have been applied in various areas [5, 6, 7, 8]. However, due to the inconvenience of using commercial closed-source models and the challenges of specific deployment, as well as the goal of achieving real AI democratization, it is crucial to propose the LLMs/MLLMs-driven manipulation systems that allows all researchers to access complete information. Based on this, manipulation systems leveraging open-source LLMs have also been proposed. However, unlike the simultaneous improvement in visual perception and instruction semantic understanding brought by incorporating commercial large model API, open-source LLMs-driven systems still rely on traditional vision models, such as YOLO [19] and Mask R-CNN [20], for the visual perception module to acquire visual modality information. Traditional vision models are typically designed for specific, single-vision tasks, which result in a lack of generality and scalability when applied to different zero-shot tasks. During the training process, these models learn from manually annotated, simple vision-label information pairs, which limits their ability to develop a more comprehensive and complex understanding of visual information when combined with rich language descriptions. This would incur the following limitations for LLMs-driven manipulation systems in open-vocabulary scenarios: (1) inability to perceive and handle unseen or zero-shot novel objects and environments; (2) inability to understand complex semantic instructions, which prevents the system from responding to more sophisticated tasks. To overcome the aforementioned limits, building upon the design principles of MLLMs, the end-to-end Vision-Language-Action Models (VLAMs) have been proposed.
@@
As for the physical security threats to AI-driven manipulation systems, [21] and [22] investigate the potential safety issues associated with using imitation learning. [23, 24, 25] explore how commercial API and open-source LLMs-driven manipulation systems experience performance degradation due to hallucination and jailbreak issues. As for the MLLMs-driven end-to-end VLAMs, no related physical security validation work has been introduced so far. Therefore, this paper conducts a comprehensive robustness and safety evaluation of open-source VLAMs, represented by LLaRA [26] and OpenVLA [27]. For the specific evaluation approaches that may arise when applying VLAMs under fully open-vocabulary scenarios within the physical world. We do not include attack methods targeting the language instructions in our threat considerations. The specific reason is that in physical attack scenarios, VLAMs mainly face relatively unchanged visual modality information, along with various semantic and linguistic template language instructions to process. Therefore, compared to constantly modifying language modality input, attacks targeting the visual modality input would have a more profound impact on the physical world. Based on this, we propose the Physical Vulnerability Evaluation Pipeline (PVEP) to assess the performance variations of VLAMs when confronting visual modality safety threats. PVEP includes most currently possible visual attacks in the physical world, such as Out-of-Distribution (OOD), Typography-based Visual Prompt (VP), and Adversarial Patch (AdvP) Attacks, to the best of our knowledge. Measured by the failure rate and the number of timesteps to complete tasks, Figure 1 shows the largest performance degradation and corresponding time delay of VLAMs before (Clean) and after suffering various physical attacks (OOD, VP, AdvP). Furthermore, through comprehensive evaluations in PVEP, we present four generalizable Analyses for VLAMs in open-vocabulary scenarios:
@@ II Background
LLMs/MLLMs-driven Manipulation systems: The recent emergence of commercial closed-source large-model APIs, represented by GPT-4 [28] and Claude [29], marks a significant step forward in the research toward Artificial General Intelligence (AGI). In contrast to closed-source models, numerous open-source LLMs and MLLMs are proposed to promote the democratization of AI development. For open-source LLMs, Flamingo [14] LLama [15], Vicuna [30] and others [31, rawte2023survey, 32] obtain good performance in various zero-shot natural language processing tasks. Open-source MLLMs, such as LLaVA, MiniGPT-4 and others  [16, 17, 18, 33], could be obtained through fusing open-source pre-trained vision encoders [34] with different open-source LLMs. Based on this progress, there are lots of commercial API-driven manipulation systems have been adopted widespread applications in different areas, such as home-services robot [5, 6], agriculture robot [7], scientific experiments robot [8]. Then, referring to the advancement of open-source LLMs, some LLMs-driven manipulation systems, such as VIMA and others [35, 36], are also proposed. Additionally, about MLLMs-driven manipulation systems, a promising trend has emerged in the form of Vision-Language-Action models (VLAMs), which involve fine-tuning large pre-trained MLLMs for robot action prediction. These models are characterized by their fusion of robot control actions directly into MLLMs backbones. RT-2 [37] and PaLM [38, 39] have garnered significant attention due to their claims of achieving promising performance across various manipulation tasks. However, both of them are closed-source. Therefore, LLaRA [26] and OpenVLA [27] as the totally open-source VLAMs are proposed. Since open-source models can be continuously evolved due to their easy accessibility, these two VLAMs may potentially revolutionize autonomous systems and human-robot interaction paradigms in the future.
@@
Safety Concerns: Different types of physical visual attacks could significantly impact the performance of various AI models. OOD attacks [40, 41, 42, 43] can consistently affect the stability of AI models. For the recently emerged LLMs and MLLMs, there are also issues such as jailbreak [44, 45, 43, 46] and hallucination [31, rawte2023survey, 47, 48] that can undermine the reliability of the final language output. Besides, for MLLMs, typography-based visual prompts [49, 50] could distract the semantics of final language output by adding simple pixel-level text to visual modality input. In addition, various kinds of AI models, including LLMs and MLLMs, as victim models, have been certified to possess adversarial vulnerabilities [51, 52, 53, 54]. The output of AI models can be altered by adversarial perturbations, leading to either deviation from the correct output (untargeted) or convergence towards a predefined incorrect output (targeted). Furthermore, the generating process of such perturbations could be classified into black and white box, depending on the amount of information attackers possess about the victim model. About the particular vulnerability of AI-driven manipulation systems, [21] and [22] investigate the potential safety issues of using diffusion policy during executing manipulation tasks. [25] explores the jailbreak threats in commercial API-driven manipulation systems. And [24] introduces the MMRo benchmark, designed to evaluate API-driven systems across four key areas: perception, task planning, visual reasoning, and safety. [23] evaluates safety challenges of open-source LLMs-driven manipulation systems in vision-language modality. However, there is no physical threat evaluation work targeting open-source VLAMs so far.
@@ III-B Threat Models
Additionally, it is crucial to analyze the impact of specific types of input image-language modality pairs on the adversarial transferability performance of VLAMs. Algorihtm 1 presents the execution process of the adversarial transfer attack from MLLMs to VLAMs. For black-box Attackers, compared to robotic vision-instruction language pairs $\{x_{rv},t_{vi}\}$ for VLAMs, which typically requires simulation or real-world scene recording to obtain, the general VQA image-prompt pairs $\{x_{gi},t_{gp}\}$ of MLLMs are often more easily available as the online dataset, such as DAQUAR [55], TallyQA[56], A-OKVQA [57] and others.
@@
[…]
@@ IV Experiments
Failure Rate (%) Clean Blur GN BC(B) BC(D) TW1 TW2 TW3 TW4 TN1 TN2 TN3 $B$ RBB GB WB
@@ IV-A Models and Simulators
For VLAMs, we adopt LLaRA [26] and OpenVLA [27] as victim models. The LLaRA and OpenVLA are evaluated on the VIMA [35] and SimplerEnv [58] simulator respectively.
@@ IV-B3 Adversarial Patch Attack
In particular, for the BB attack, we utilize minimal information about the victim VLAMs. We employe only the victim VLAMs’ base model (pre-fine-tuning) as a surrogate model. For instance, for the victim model LLaRA, we use its base model LLaVA as the surrogate model. The adversarial patch was trained using 5000 images from ImageNet and 200 general Visual Question Answering (VQA) prompts selected from [59]. For the RBB attack, based on BB, we provide additional data information for training the adversarial patch. We collect an additional 2000 Images and 200 prompts of robotics manipulation scenarios when using victim VLAMs for inference. For the GB attack, we use the victim VLAMs and the general VQA images/prompts in BB for adversarial patch training. As for the WB attack, we directly use the victim VLAMs and all available images and prompts from both BB and RBB. By systematically comparing these four levels of adversarial patch attacks, we can assess the minimum level of model and data information required to generate an effective adversarial patch against VLAMs.
@@
Failure Rate (%) Clean Blur GN BC(B) BC(D) TW1 TW2 TW3 TW4 TN1 TN2 TN3 $B$ RBB WB